16 May 2011

Fine weather continues - we make good progress

Sarah at work in the kitchen, preparing yet another meal

As weather continues to be good to us, work on getting camp into shape before reaching full manning this week is going fine. CFA setup in the laboratory is progressing and the drillers are trying out a new experimental drill head. As with all tests, there is a troubleshooting phase and the drillers have been working at a slower pace to see where the problems are. The new head makes perfect ice cores, but the ice chips produced during drilling are not moved from the head to the chip chamber efficiently enough. For the planned 420 m drilling this is not a problem at all, as the drillers can switch back to our well tested normal drill head and proceed with that. Tests like this are necessary, if we want to advance our technology.

There is one item on the “what we have done today” list that is always missing (except on Saturdays), and that is Sarah providing food for all in camp, every day. Not only does she provide us with meals every day, she also plans food purchases, keeps check on food
inventory and makes sure we have a balanced diet.

What we have done today:

1.  Drilling and logging of ice cores.
2.  Removing snow and leveling surface in the 2nd weather port area.
3.  Excavating last two pallets with fuel and drilling fluid from last year.
4.  Excavating and pulling all three fuel tanks into position near the apron.
5.  Zigzag grooming the skiway.
6.  Raised and reset all approach markers.
7.  Setting up CFA system.
8.  Again today, Sarah provided us with two nice meals.

Drillers report: Drilling today was slowed due to issues with poor chip transport. Most likely this is caused by the new experimental drill head. Drillers are investigating the problems. Drillers’ depth: 25.5 meter.

Weather: Clear. Temp. - 15 °C  to - 22 °C , 12 knots from SSE. Unrestricted visibility.

FL, J.P. Steffensen 

← Previous entry   Next entry →